If a rocket is launched from Houston, Texas, with an initial velocity of 1600 feet per second, when will the rocket be 14,400 feet high?
Answer by stanbon(75887)

If a rocket is launched from Houston, Texas, with an initial velocity of 1600 feet per second, when will the rocket be 14,400 feet high?
---------------
h(t) = -16t^2+1000t
------
Solve:: -16t^2+1000t-14,400 = 0
(2x-45)(-8t+40) = 0
---------------
Positive solution:
x = 22.5 seconds
-----------------------------
Cheers,
Stan H.
-----------------